OpenStudy (anonymous):

This equation represents the population of Carrisville in y years from now. Substitute 5 years for y and find the population for 5 years from now. Round your answer to the nearest whole number. A. 2,760 B. 12,750 C. 2,550 D. 12,500

OpenStudy (alfie):

You should give us the equation in order to plug in 5years for y :P

OpenStudy (anonymous):

sorry p=2500(1.02)y

OpenStudy (anonymous):

12750.

OpenStudy (alfie):

Just plug in "5" instead of y and do the products. It comes B 12750

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

I think it was meant to be \[p = 2500 \times 1.02^{y}\]

OpenStudy (anonymous):

It seems like an exponential growth question and why would give anyone a formula like p = 2500(1.02)y